A 1000 kg car is travelling on the highway at a constant speed of 110 km/h.



1
Expert's answer
2021-03-31T09:32:03-0400

So,


A 1000 kg car is travelling on the highway at a constant speed of 110 km/h. The highway makes a turn that is banked at an angle of 10° and has a radius of 160 m. What is the coefficient of friction between the car tires and the pavement in the curve?



"F_{net}=m\cdot\frac{v^2}{r}"


"F_{net}=N\sin10°+\mu N\cos10°"



"mg+\mu N\sin10°=N\cos10°"


"N\sin10°+\mu N\cos10°=m\cdot\frac{v^2}{r}"


We have


"\mu=\frac{(v^2/(gr))\cos10°-\sin10°}{(v^2/(gr))\sin10°+\cos10°}=\frac{(30.6^2/(9.8\cdot160)\cos10°-\sin10°}{(30.6^2/(9.8\cdot160)\sin10°+\cos10°}=0.381" . Answer